While skydiving, your parachute opens and you slow down from 50.0 m/s to 8.00 m/s in 0.800 s. (a) Determine the distance you fall while the parachute is opening (b) Determine the magnitude of your acceleration during this time interval (c) Divide the acceleration found in (b) by g-9.80 m/s. Typically, some people feel faint if this exceeds 5 (5 gs). Will you feel faint? (d) What assumptions did you make in solving this problem?
Answer #1

Given,

u = 50 m/s ; v = 8 m/s ; t = 0.8 s

a)We know from eqn of motion

v^2 = u^2 + 2 a S

S = (v^2 - u^2)/2 a

S = (8^2 - 50^2)/2 x -9.81 = 124.16 m

Hence, S = 124.16 m

b)v = u + at

a = (v - u)/t

a = (8 - 50)/0.8 = -52.5 m/s^2

Hence, a = -52.5 m/s^2(negative as its deceleration)

c)a = 52.5/9.81 x g = 5.35 g

As the value of acceleration is exceeding 5g, yes you will feel faint.

d)We assumed that its freely falling without air drag.